# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

### v2

@@ Abstract
[…]
@@ 1 Introduction
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [5, 28, 46, 42, 4] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. As a result, the nature of video redundancy remains underexplored. In this work, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
@@
To maintain visual context, redundant tokens within each segment must be effectively merged. A state-of-the-art method VisionZip [42] applies uniform token sampling followed by merging redundant tokens [4] as shown in Figure 1(b1). However, its token selection strategy is content-agnostic, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1(b2). Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [26]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
@@
[…]
The main contributions are summarized as follows: (1) We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity. (2) We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details. (3) Our FastVID achieves state-of-the-art performance across diverse video benchmarks and maintains robust accuracy even under extreme compression.
@@ 2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 7, 34] and MLLMs [35, 21, 22, 17, 3, 23], there has been growing interest in Video LLMs. These models can be categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
@@
Video LLMs with training-time compression [31, 24, 40, 15, 30, 49] aim to significantly reduce the number of video tokens, enabling longer video sequences. PLLaVA [40] introduces an adaptive average structure pooling to extend image LLMs. Video-ChatGPT [24] extracts both spatial and temporal features through temporal and spatial pooling respectively. Chat-UniVi [15] progressively clusters visual tokens and provides multi-scale features. LongVU [30] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [49] explores scaling consistency and uses the Perceiver Resampler [14].
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision being widely adopted due to its adaptability and superior performance. Therefore, we focus on inference-time acceleration for general Video LLMs. To evaluate our method, we select LLaVA-OneVision and LLaVA-Video as representative models.
@@
Recent studies [5, 28, 46, 42, 4, 39] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [5] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [28] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [46] proposes a token recycling strategy to aggregate and reconstruct tokens to be pruned. VisionZip [42] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [12, 33, 13] have gained increasing attention. FrameFusion [12] applies both merging and pruning across successive shallow LLM layers, but repeated pruning operations adversely affect overall efficiency. DyCoke [33] merges tokens across frames and applies dynamic KV cache reduction. However, its pruning during the prefilling stage struggles to achieve substantial token reduction while maintaining accuracy. PruneVID [13] clusters video tokens and selects those most relevant to query tokens, but this dependency on clustering introduces significant latency during compression. Despite these advances, efficient and accurate pruning under large token reduction remains unsolved.
In this paper, we focus on preserving temporal and visual context for spatiotemporal token pruning at the prefilling stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [8], KV cache, and a plug-and-play design for seamless integration into existing Video LLMs. Our method achieves robust performance even under extreme compression rates, offering a practical solution for fast Video LLMs.
@@ 3.1 FastVID
Video LLMs face significant inference challenges due to the large number of frames in a video and the numerous vision tokens within each frame. Typically, the total token count often exceeds the LLM’s context length, making direct processing computationally infeasible. Although uniform frame sampling and token pooling methods are commonly employed to reduce complexity, they still leave a substantial number of redundant video tokens (e.g., $32\times 196=6272$ video tokens in LLaVA-OneVision [16]).
To address these limitations, we propose FastVID, a novel method that accelerates inference by reducing redundant tokens while preserving critical temporal and visual context. First, we introduce Dynamic Temporal Segmentation (DySeg) in Section 3.2, which partitions a video into temporally ordered segments with high redundancy. Within each segment, we introduce Density Spatiotemporal Pruning (STPrune) in Section 3.3, which retains essential visual context and salient details.
@@
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. Specifically, DySeg dynamically partitions video tokens into temporally ordered, high-redundancy segments, while STPrune performs density-based pruning within each segment, enhancing efficiency without loss of key information. The final retained video tokens, combined with query tokens, is then fed into the LLM to generate the response.
@@ 3.2 Dynamic Temporal Segmentation
To enable effective pruning, DySeg promotes high spatiotemporal redundancy within each segment by minimizing similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
@@
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the $F-1$ transition similarities for $F$ sampled frames. To achieve fine-grained segmentation, we select transitions that satisfy the following conditions:
@@
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the $c-1$ least similar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. Each transition in the union $\mathbf{S}$ marks a boundary between segments. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3(d), our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
@@ 3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for segment visual context and Attention-based Token Selection (ATS) for salient details. For a segment of $P$ frames, we retain $rPN$ tokens in total, where $N$ is the number of tokens per frame, and $r$ denotes the retention ratio. These tokens are generated by DTM and ATS. Specifically, $drFN$ tokens are generated by DTM and $(1-d)rFN$ tokens are generated by ATS, where $d$ controls the token allocation between the two modules.
@@
To further clarify DTM, Figure 4(b) illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [26, 10] to compute the density score. For each token in the anchor frame $[v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
@@
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while remaining relatively distant from other high-density tokens. This selection ensures that anchor tokens are both representative and distinctive. Next, for each anchor frame, Similarity-based Assignment assigns each token in the segment to the nearest anchor token using cosine similarity. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $[b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\alpha a+\frac{1-\alpha}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\alpha$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [32] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [13] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding.
@@
Figure 4: Comparison of DTM and Cluster-based Token Merging. (a) Cluster-based methods aggregate tokens within each cluster and concatenate them, leading to a loss of positional information and disruption of the spatiotemporal structure in video data. (b) Our DTM updates anchor tokens while preserving their original positional information, maintaining structural coherence. Please refer to Table 5 for a detailed quantitative comparison.
@@
Specifically, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top $(1-d)rN$ tokens with the highest attention scores to preserve critical visual details.
@@ 4.1 Experimental Settings
Table 1: Comparison of state-of-the-art methods on LLaVA-OneVision [16]. The A%/Bretention ratio indicates that Aof the LLM input tokens are retained, and subsequently compressed to Bduring the LLM forward pass. The best performance among those with similar retention ratios $R$ is highlighted in bold. TFLOPs related to video tokens are reported (see the appendix for details).
Method Retention Ratio $R$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
@@
DyCoke [33]${}_{\text{CVPR'25}}$ 32.5% 14.13 56.3 56.6 62.1 57.1 68.1 56.7 46.7 58.0 97.8
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 13.45 54.7 55.5 61.5 56.2 68.0 54.6 46.0 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 53.7 51.2 58.5 54.1 61.6 53.4 47.2 54.4 91.7
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 56.6 55.7 64.8 58.0 68.6 57.7 47.7 58.8 99.1
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 10.73 49.5 48.1 55.8 51.0 61.1 48.6 43.2 51.1 86.2
@@
FastV [5]${}_{\text{ECCV'24}}$ 100%/20% 11.38 54.1 56.6 61.2 56.2 66.8 54.6 47.2 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 53.0 50.0 57.1 53.0 60.8 51.0 47.1 53.3 90.0
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 55.8 55.4 64.2 58.0 68.6 57.0 48.3 58.4 98.5
@@
FastV [5]${}_{\text{ECCV'24}}$ 100%/15% 9.35 53.2 54.9 59.8 54.7 65.1 53.4 45.7 55.7 93.9
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 50.3 46.9 54.4 49.5 55.8 49.3 43.3 50.3 84.8
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 54.3 53.9 63.1 55.5 63.0 54.4 49.1 56.7 95.6
@@
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 7.36 51.7 52.1 57.7 52.4 60.9 51.4 45.0 53.5 90.2
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 44.4 43.5 51.5 46.0 50.4 45.8 41.8 46.4 78.3
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 51.7 48.3 59.7 52.8 59.4 52.0 46.9 53.1 89.6
PruneVID [13]${}_{\text{ACL'25}}$ 10.1% 4.23 54.2 53.8 62.3 55.9 66.4 52.9 48.3 56.6 95.4
@@
Method Retention Ratio $R$ Newline Tokens $M$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
@@
DyCoke [33]${}_{\text{CVPR'25}}$ 32.1% 256 27.1 59.3 57.9 65.7 61.6 74.6 51.1 61.1 96.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 832/568.7 29.2 58.0 58.3 63.9 61.0 71.3 51.0 60.3 94.8
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 56.4 54.1 62.1 58.6 66.3 51.2 57.8 90.9
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 58.3 58.3 66.6 61.9 73.3 52.2 61.5 96.7
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 208 20.7 50.8 53.0 56.9 56.1 65.8 48.9 54.2 85.2
@@
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 832/311.8 16.2 55.8 55.4 58.9 57.9 67.6 48.6 57.0 89.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 46.3 46.6 52.2 49.5 54.2 44.3 48.7 76.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 56.6 53.6 61.7 58.7 67.6 50.1 57.7 90.7
@@
Table 3: Efficiency Comparison on LLaVA-OneVision [16]. The prefilling time, defined as the latency to the first generated token, is measured on VideoMME using an A100 GPU.
Method # Token TFLOPs Prefilling Time (ms) Avg. Acc.
@@
Vanilla 6272 (100%) 48.82 (100%) 476.3 476.3 (1.0$\times$) 59.3 (100%)
PruneVID [13] 635.6 (10.1%) 4.23 (8.7%) 5.2 32.0 64.3 101.5 (4.7$\times$) 56.6 (95.4%)
FastVID 608 (9.7%) 4.04 (8.3%) 0.5 5.6 61.1 67.2 (7.1$\times$) 58.1 (98.0%)
Implementation Details. We apply our method to two representative Video LLMs: LLaVA-OneVision-7B [16] is a general Video LLM, while LLaVA-Video-7B [47] employs newline tokens based on positional information of video tokens. These models encompass different architectural characteristics, ensuring a robust evaluation of our approach. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\alpha=0.6$ for all experiments. For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM. We test with $r\in\{25\%,20\%,15\%,10\%\}$. For LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. We test with $r\in\{25\%,10\%\}$. We conduct all evaluations using LMMs-Eval [44] on A100 GPUs.
Compared Baselines. (1) For image compression, we adopt both the widely used classic method FastV [5] and the current state-of-the-art approach VisionZip [42]. VisionZip prunes tokens in the last layer of the vision encoder, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling. (2) For video compression, we compare two recent methods, DyCoke [33] and PruneVID [13]. Our FastVID focuses on pruning during the prefilling stage. To ensure fairness, we reimplement these baselines without pruning in the decoding stage. PruneVID applies two-stage pruning on both input tokens and the LLM’s 10th layer. PruneVID refers to the variant that prunes only input tokens. More implementation details of these baselines are provided in the appendix.

### v3

@@ Abstract
Video Large Language Models have demonstrated strong video understanding capabilities, yet their practical deployment is hindered by substantial inference costs caused by redundant video tokens. Existing pruning techniques fail to effectively exploit the spatiotemporal redundancy present in video data. To bridge this gap, we perform a systematic analysis of video redundancy from two perspectives: temporal context and visual context. Leveraging these insights, we propose Dynamic Density Pruning for Fast Video LLMs termed FastVID. Specifically, FastVID dynamically partitions videos into temporally ordered segments to preserve temporal structure and applies a density-based token pruning strategy to maintain essential spatial and temporal information. Our method significantly reduces computational overhead while maintaining temporal and visual integrity. Extensive evaluations show that FastVID achieves state-of-the-art performance across various short- and long-video benchmarks on leading Video LLMs, including LLaVA-OneVision, LLaVA-Video, Qwen2-VL, and Qwen2.5-VL. Notably, on LLaVA-OneVision-7B, FastVID effectively prunes 90.3% of video tokens, reduces FLOPs to 8.3%, and accelerates the LLM prefill stage by $\textbf{7.1}\times$, while maintaining 98.0% of the original accuracy. The code is available at https://github.com/LunarShen/FastVID.
@@ 1 Introduction
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [6, 32, 56, 51, 5] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. As a result, spatiotemporal redundancy remain insufficiently explored. In this work, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
@@
Given high-redundancy segments, we aim to preserve visual context by effectively consolidating spatial and temporal information within each segment. A state-of-the-art method VisionZip [51] applies uniform token sampling followed by merging redundant tokens [5] as shown in Figure 1(b1). However, its token selection strategy is content-agnostic, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1(b2). Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [30]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
@@
To evaluate the generalization capability of our approach, we evaluate it on leading Video LLMs, LLaVA-OneVision [19], LLaVA-Video [57], Qwen2-VL [45], and Qwen2.5-VL [4]. To further validate its effectiveness, we perform extensive experiments on MVBench [21], LongVideoBench [46], MLVU [58], and VideoMME [12]. These benchmarks cover a wide range of video complexities and durations, ensuring a comprehensive evaluation. Notably, on LLaVA-OneVision-7B, FastVID effectively prunes 90.3% of video tokens, reduces FLOPs to 8.3%, and accelerates the LLM prefill stage by $\textbf{7.1}\times$, while maintaining 98.0% of the original accuracy across all benchmarks.
The main contributions are summarized as follows: (1) We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity. (2) We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details. (3) Our FastVID achieves state-of-the-art performance across diverse Video LLMs and benchmarks, while maintaining robust accuracy even under extreme compression.
@@ 2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 8, 42] and MLLMs [43, 25, 26, 20, 3, 27, 15, 14], there has been growing interest in Video LLMs. These models can be categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
@@
Video LLMs with training-time compression [38, 28, 48, 18, 37, 60] aim to significantly reduce the number of video tokens, enabling longer video sequences. Chat-UniVi [18] progressively clusters visual tokens and provides multi-scale features. LongVU [37] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [60] explores scaling consistency and uses the Perceiver Resampler [17].
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision and the Qwen-VL series being widely adopted due to its adaptability and superior performance. Therefore, we focus on inference-time acceleration for general Video LLMs.
@@
Recent studies [6, 32, 56, 51, 5, 47, 23, 59] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [6] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [32] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [56] proposes a token recycling strategy to aggregate and reconstruct tokens. VisionZip [51] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [13, 41, 16, 36, 34, 40] have gained increasing attention. FrameFusion [13] applies both merging and pruning across successive shallow LLM layers, but repeated pruning operations adversely affect overall efficiency. DyCoke [41] merges tokens across frames and applies dynamic KV cache reduction. However, its pruning during the prefill stage struggles to achieve substantial token reduction while maintaining accuracy. PruneVID [16] clusters video tokens and selects those most relevant to query tokens, but this dependency on clustering introduces significant latency during compression. Despite these advances, efficient and accurate pruning under large token reduction remains unsolved.
@@ 3.1 FastVID
In this paper, we focus on preserving temporal and visual context at the prefill stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [9], KV cache, multi-turn conversations, and a plug-and-play design for seamless integration into existing Video LLMs. Our method achieves robust performance even under extreme compression rates, offering a practical solution for fast Video LLMs.
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. Specifically, Dynamic Temporal Segmentation (DySeg) in Section 3.2 dynamically partitions video tokens into temporally ordered, high-redundancy segments, while Density Spatiotemporal Pruning (STPrune) in Section 3.3 performs density-based pruning within each segment, enhancing efficiency with minimal loss of key information. The final retained video tokens, combined with query tokens, is then fed into the LLM to generate the response.
@@ 3.2 Dynamic Temporal Segmentation
To enable effective pruning, DySeg induces high spatiotemporal redundancy within each segment by minimizing similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
@@
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the F1F-1 transition similarities for $F$ sampled frames. To achieve dynamic segmentation, we select transitions that satisfy the following conditions:
@@
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the c1c-1 most dissimilar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. Each transition in the union $\mathbf{S}$ marks a boundary between segments. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3(d), our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
Figure 4: Comparison of our proposed DTM and Cluster-based Token Merging. (a) Cluster-based methods aggregate tokens within each cluster and concatenate them, leading to a loss of positional information and disruption of the spatiotemporal structure in video data. (b) Our DTM updates anchor tokens while preserving their original positional information, maintaining structural coherence. Please refer to Table 7 for a detailed quantitative comparison.
@@ 3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for segment visual context and Attention-based Token Selection (ATS) for salient details. For a segment of $P$ frames, we retain $rPN$ tokens in total, where $N$ is the number of tokens per frame, and $r$ denotes the retention ratio. These tokens are generated by DTM and ATS. Specifically, $drPN$ tokens are generated by DTM and (1d)rPN(1-d)rPN tokens are generated by ATS, where $d$ controls the token allocation between the two modules.
@@
To further clarify DTM, Figure 4(b) illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [30, 11] to compute the density score. For each token in the anchor frame $v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
@@
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while remaining relatively distant from other high-density tokens. This selection ensures that anchor tokens are both representative and distinctive. Next, for each anchor frame, Similarity-based Assignment assigns each token in the segment to the nearest anchor token using cosine similarity. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\beta a+\frac{1-\beta}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\beta$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [39] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [16] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding.
@@
Specifically, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top (1d)rN(1-d)rN tokens with the highest attention scores to preserve critical visual details.
@@ 4.1 Experimental Settings
Table 1: Comparison of state-of-the-art methods on LLaVA-OneVision [19]. The A%/B% retention ratio indicates that A% of the LLM input tokens are retained, and subsequently compressed to B% during the LLM forward pass. The best performance among those with similar retention ratios $R$ is highlighted in bold. TFLOPs related to video tokens are reported (see Appendix A for details).
Method
Retention
Ratio $R$
TFLOPs
MVBench
LongVideo
Bench
MLVU VideoMME Avg. Acc.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Method
Retention
Ratio $R$
# Newline
Tokens $M$
TFLOPs
MVBench
LongVideo
Bench
MLVU VideoMME Avg. Acc.
@@
[…]
@@
[…]
@@
Table 3: Comparison of state-of-the-art methods on Qwen2-VL [45]. We set the maximum number of video tokens fed into the LLM to 16384, and the maximum number of sampled frames to 768. FastV performs pruning based on LLM attention scores, but its eager-attention implementation materializes the full attention matrix in memory, leading to OOM errors due to the large number of video tokens.
Method # Token TFLOPs VideoMME
Short Medium Long Overall
Vanilla 13447.1 100% 124.0 100% 74.1 60.4 54.3 63.0 100%
[…]
FastVID 3349.3 24.9% 24.1 19.4% 72.7 58.3 50.7 60.6 96.2%
Table 4: Comparison of state-of-the-art methods on Qwen2.5-VL [4]. We set the maximum number of video tokens fed into the LLM to 16384, and the maximum number of sampled frames to 768.
Method # Token TFLOPs VideoMME
Short Medium Long Overall
Vanilla 13447.1 100% 124.0 100% 76.7 68.2 56.9 67.3 100%
[…]
FastVID 3240.7 24.1% 23.2 18.7% 74.3 61.3 52.8 62.8 93.3%
Table 5: Efficiency Comparison on LLaVA-OneVision [19]. The prefill time, defined as the latency to the first generated token, is measured on VideoMME using an A100 GPU.
Method # Token TFLOPs Prefill Time (ms) Avg. Acc.
@@
[…]
Implementation Details. We apply our method to leading Video LLMs: LLaVA-OneVision [19], LLaVA-Video [57], Qwen2-VL [45], and Qwen2.5-VL [4]. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\beta=0.6$ for all experiments. For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM. We experiment with $r\in\{25\%,20\%,15\%,10\%\}$. For LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. We experiment with $r\in\{25\%,10\%\}$. For the Qwen-VL series, which samples up to 768 frames, we discard highly redundant frames and set $\tau$ to its optimal value. We conduct all evaluations using LMMs-Eval [54] on A100 GPUs.
Compared Baselines. (1) For image compression, we adopt both the widely used classic method FastV [6] and the current state-of-the-art method VisionZip [51]. VisionZip prunes tokens in the last layer of the MLLM’s ViT, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling. (2) For video compression, we compare two recent methods, DyCoke [41] and PruneVID [16]. Our FastVID focuses on pruning during the prefill stage. To ensure fairness, we reimplement these baselines without pruning in the decode stage. PruneVID applies two-stage pruning on both input tokens and the LLM’s 10th layer. PruneVID refers to the variant that prunes only input tokens. More implementation details of these baselines are provided in Appendix A.